Solve by system of equations.
THe sum of the digits of a two digit number is 13. If the digits of the number are reversed,the new number is 27 less than the origional number. Find the number.
i have so far
10y+x = 10x+y-27 ??
show all work please

Answer by mananth(16946)


You have not set up another equation
let the number be x y
x+y= 13.....................1
10y+x = 10x+y-27
10y-y+x-10x=-27
9y-9x==-27
/9
y-x=-3......................2
..
add 1 & 2
2y= 13-3
2y=10
/2
y=5
..
x+y = 13
so x=8
..
The number is 85
..
m.ananth@hotmail.ca